Item 2.9

A 9.9-V battery is connected in series with a 36-mH inductor, a 180-Ω resistor, and an open switch.

Part A: What is the current in the circuit 0.120 ms after the switch is closed?

Part B: How much energy is stored in the inductor at this time?
Answer #1

For this particular question we will use the formula given below-

Now at t = 0.120 ms , current will be -

iL = ( 9.9 * (1-e-0.120*10-3 / 0.2 *10-3 )) / 180

iL = ( 9.9 * (1-e-0.6 )) / 180

iL = 0.024815 A = 24.815 mA

Part B)-

Energy in inductor = 0.5*L*iL2

E = 0.5 * 36*10-3 * 0.000616

E = 0.000011 J = 11 * 10-6 J

Thankyou !!! Have Fun

Please Rate High.